Express the real part of the following signal in the form Ae-atcosωt+θ where A, a, ω...

  1. Express the real part of the following signal in the form Ae-atcosωt+θ where A, a, ω and θ are real and -π<θ≤π
    1. xt= e-5tsin⁡6t-π4 (4 points)

  1. xt= 9je-6+j400t   (8 points)
Answer #1

-St ca) X(t) = è st sinlet - To(43) Re{z(t)} = Reşist sinkt - TT (4))} Splitting -40 +(-9%) = Reş -st e sin (6t + 1/2 9%)} =2) x(+) = 9; et 6+ juoolt Re{x(t)} Re{aje 6t4 400t } 30 Re { qj ebt juoot } = (40+ isino -bt - Re{qje [cos (400 t) + jsinyoo